OpenStudy (anonymous):

Solve: -4(2x - 1) = -6(x + 2) - 2

OpenStudy (anonymous):

-4(2x - 1) = -6(x + 2) - 2 -8x + 4 = -6x - 12 - 2 -8x + 4 = -6x - 14 4 + 14 = -6x + 8x 18 = 2x 18 / 2 = x x = 9